OpenStudy (anonymous):

if x is 20% of y and y is 40% of z, what is x in terms of z?

OpenStudy (lgbasallote):

x is 20% of y so \[x = \frac{20}{100} \times y\] y is 40 % of z so \[y = \frac{40}{100} \times z\] so that means \[x = \frac{20}{100} \times \frac{40}{100} \times z\] ^i just substituted y does that help?

OpenStudy (lgbasallote):

@kellyXD are you there?

OpenStudy (anonymous):

oh okay thanks

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):